Find the general solution of the differential equations taking into account the initial conditions using the parameter variation method :

. y + 4y = t y(0) = y(0) = 0 et y(0) = 13 y – 3y + 2y = ttet ; y(0) = 1; y(0) = Let y(0) 2
Answer #1

And I y + 4y =t t (D3 +4D)Y where D = d dt A. E m3 + 4m 0 mlm² + 4) = 0 m +21 -21 a eot + ( Cus(at) + C Sinlat) C + 2 Cos(at)Cosat 0 - 2sin (24) 10 0 - asinat 1 - 4 Cusat = Now dt t² 8 sta ses - StoSet Uy = - to [ 24 Sin (at) + cos(at)] t - 2 Cosat 8yo 61 - 24 Sin(at) + 20, Cos(at) + + Y - 4 G Cos(at) - 4G Sin (at) + 1 + Given yllo) > G - 2₂ xo + 2Gx1 to Also y (o) = 1 -Let y = 1 Y = et Y = et W= 1 et ezt 0 et ge3t Zet 4e2t O et Wia 0 et ezt O 2e4 E et et et 1 4 eet 1 o O 0 ezt 2 ezt 4 et -+2 + tla +7 - tet 4 Hence C.F. + P.I. 11 7- G + Get + geht + 4 + 2 + + 7 3 t t -tet 8 y= Y Get +26e²t tt + 3 -tet - et (21 Y